# MAPLE: Self-supervised Learning-Enhanced Nonlinear Dimensionality Reduction for Visual Analysis

### v1

@@ Abstract
We present a new nonlinear dimensionality reduction method, MAPLE, that enhances UMAP by improving manifold modeling. MAPLE employs a self-supervised learning approach to more efficiently encode low-dimensional manifold geometry. Central to this approach are maximum manifold capacity representations (MMCRs), which help untangle complex manifolds by compressing variances among locally similar data points while amplifying variance among dissimilar data points. This design is particularly effective for high-dimensional data with substantial intra-cluster variance and curved manifold structures, such as biological or image data. Our qualitative and quantitative evaluations demonstrate that MAPLE can produce clearer visual cluster separations and finer subcluster resolution than UMAP while maintaining comparable computational cost.
@@ I Introduction
High-dimensional datasets are rapidly growing across scientific domains, from single-cell transcriptomics to large-scale image collections. Advances in genomics, machine learning, and related fields often produce measurements with tens of thousands of features, such as genes, pixels, or descriptors, that naturally place the data in very high dimensions. As humans cannot directly see high-dimensional spaces, effective visualization is essential to analyze these datasets. Dimensionality reduction (DR) methods make high-dimensional data visually explorable by projecting them into a low-dimensional space, typically 2D or 3D. Successful DR results can facilitate data annotation, reveal rare populations, and lead to the generation of new hypotheses.
@@
A wide range of DR methods has been proposed for visualization [espadoto2019toward, jeon2025unveiling]. Linear DR methods, including principal component analysis (PCA) [hotelling1933analysis], remain popular for their interpretability [cunningham2015linear], but are limited in capturing complex, curved data structures. Nonlinear DR methods address this limitation by prioritizing local neighborhood relationships, allowing them to better represent data lying on complex manifolds [van2009dimensionality]. Among these, uniform manifold approximation and projection (UMAP) [mcinnes_umap_2020, healy2024umap] has become a method of choice across many domains, alongside alternatives such as $t$-SNE [maaten2008visualizing] and PHATE [moon_PHATE_2019]. UMAP is valued for its scalability, graph-based formulation, and ability to capture nonlinear structure without strong parametric assumptions. UMAP constructs a weighted $k$-nearest neighbor (kkNN) graph in the input space to model an underlying manifold and then optimizes a low-dimensional layout so that the structure of the layout aligns with the graph according to a cross-entropy objective. This procedure is computationally efficient, and its cross-entropy objective often yields layouts with clear cluster separation.
@@
The quality of the resulting UMAP layout is strongly influenced by the quality of the weighted kkNN graph. In practice, this graph may include connections that bring semantically distinct points close together in the low-dimensional layout, and vice versa. We argue that a key limitation of UMAP lies not in the kkNN algorithm itself—which can serve as an effective proxy for nonlinear structure—but in how the graph is constructed, specifically the choice of data representation and distance space used to define neighborhood relationships. More specifically, standard kkNN construction based on raw high-dimensional data and predefined distance metrics faces the following two issues. First, as dimensionality increases, pairwise distances in high-dimensional space (e.g., Euclidean distances) tend to concentrate toward similar values [franccois2007concentration], making kkNN graphs increasingly sensitive to noise. Second, many datasets lie on curved manifolds [kocabas2019self], where points that appear as neighbors in the kkNN graph may actually belong to distinct semantic regions [conwell2024large]. In 2D or 3D layouts, these effects result in local clusters with spurious overlaps, mixing groups that are semantically distinct. Recent theoretical work has also analyzed these issues of UMAP [damrich_umap_2021, bohm2022attraction], emphasizing its limitations in capturing subtle geometric differences.
@@
To address the above fundamental challenges, we introduce a novel nonlinear DR method, MAPLE (short for manifold-aware projection with learned edges). A central idea in MAPLE is that explicitly learning neighborhood relationships can substantially improve the quality of low-dimensional layouts. MAPLE refines neighborhood definitions by adapting multi-view self-supervised learning (MVSSL) [tsai2021selfsupervised, galvez2023role] to the context of nonlinear DR. Unlike standard methods that treat a weighted kkNN graph directly constructed from raw high-dimensional data as a fixed representation of the underlying manifold, MAPLE regards the initial neighbor relationships as weak, tentative self-supervised signals. We employ a learning phase to iteratively refine these signals. Instead of directly operating in the input high-dimensional space, we employ encoder and projector networks [grill2020bootstrap, kalantidis2021tldr, zbontar_barlow_2021] to map data into an embedding space where the distance relationships are learned rather than fixed by a predefined distance metric. A refined graph is then constructed based on these optimized distance relationships. The core of this learning phase is the maximum manifold capacity representation (MMCR) objective [yerxa2023mmcr], which compresses variance among locally similar points (i.e., neighbors) while amplifying variance across dissimilar ones (i.e., non-neighbors). The outcome of this learning phase is a refined, weighted kkNN graph that is subsequently laid out in a low-dimensional space using UMAP’s cross-entropy optimization.
@@
We evaluate MAPLE against multiple DR methods, using a set of quality measures. Additionally, we qualitatively evaluate MAPLE through visual analysis on image classification datasets. We further demonstrate MAPLE’s utility in case studies on single-cell transcriptomics in collaboration with a bioinformatics expert (a co-author of this work). Our results show that MAPLE reveals fine-grained, meaningful patterns that support deeper downstream analysis. Our contributions are as follows:
@@
Terminology. MAPLE produces two types of low-dimensional representations. To avoid ambiguity, we use the following terminology throughout the paper:
@@ II Background and Related Work
DR aims to extract low-dimensional representations from high-dimensional data while preserving meaningful structure. Nonlinear DR methods [van2009dimensionality] are designed to capture data patterns lying on nonlinear manifolds (e.g., curved surfaces). From the perspective of manifold modeling, nonlinear DR methods can be categorized by how they represent neighborhood structure. Graph-based modeling constructs a kkNN graph or a similar structure to capture local proximities along a manifold. This category includes early spectral and geodesic methods, such as ISOMAP [tenenbaum2000global] and Laplacian eigenmaps [belkin2003laplacian], as well as more recent visualization-oriented methods, including LargeVis [tang2016visualizing], $t$-SNE [maaten2008visualizing], and UMAP [mcinnes_umap_2020]. Diffusion-based modeling (e.g., spectral embeddings [meila2001spectral] and PHATE [moon_PHATE_2019]) simulates diffusion processes (e.g., random walks) on the data and uses transition probabilities as similarities to preserve manifold structure. Kernel-based modeling [mika1998kernel, gisbrecht2015parametric] employs kernel functions to implicitly map data into higher-dimensional feature spaces, enabling the modeling of complex shapes. Our new method, MAPLE, belongs to the graph-based family. We specifically aim to advance graph-based modeling, using UMAP (a widely used state-of-the-art method) as our primary baseline.
@@ II-A UMAP
where $\rho_{i}$ and $\sigma_{i}$ are local connectivity parameters automatically determined by UMAP to distribute data points uniformly over the manifold. The weight, $w_{ij}$, reflects the strength of neighborhood relationships between points, $i$ and $j$ (more precisely, membership strength of a 1-simplex between them [mcinnes_umap_2020]).
@@
$\arg\min_{\mathbf{Y}}\sum_{i\neq j}\Big[w_{ij}\log\!\left(\frac{w_{ij}}{v_{ij}}\right)-(1-w_{ij})\log\!\left(\frac{1-w_{ij}}{1-v_{ij}}\right)\Big].$ (2)
@@
As described in Eq. 1, UMAP assumes that data points are uniformly distributed on the manifold under some Riemannian metric. Under this assumption, any ball of fixed volume should contain approximately the same number of points, regardless of its position on the manifold. This motivates the use of fixed-size kkNNs to represent local relationships. In practice, however, real data are rarely uniform: they often vary in density or lie on curved manifolds. As a result, kkNNs constructed with a conventional metric (e.g., Euclidean distance) and the connectivity parameters, $\rho_{i}$ and $\sigma_{i}$, derived from them can misrepresent geodesic proximity. Since UMAP’s cross-entropy objective optimizes only over the weighted kkNN graph, any mismatch introduced during graph construction directly propagates into the final layout. This highlights a research gap: improving neighborhood construction is essential for achieving low-dimensional layouts that faithfully reflect the underlying manifold geometry.
@@ II-C Self-supervised Learning and Integration into DR
Integration of SSL into DR.  Several works have integrated SSL into DR to improve layout quality. These works share common critiques of the graph construction phase, including unreliable distances in high-dimensional space [bohm2022unsupervised], semantically unstable neighborhoods [hu2023your], and distortion caused by curvature or density differences [kalantidis2021tldr, feng2024ncldr]. SSL objectives can help learn more semantically meaningful results. For example, $t$-SimCNE [bohm2022unsupervised] combines the contrastive learning method, SimCLR, with a $t$-SNE-like loss to improve similarity estimation of data points. However, as it relies on SimCLR, $t$-SimCNE is only applicable for image datasets. TLDR [kalantidis2021tldr] employs a single nearest neighbor as the self-supervised signal and Barlow Twins [zbontar_barlow_2021] as the SSL loss. While effective for data compression and preprocessing, TLDR is not designed for visualization purposes. Moreover, its reliance on a fixed $1$-nearest neighbor graph prevents it from addressing misaligned neighborhood issues. Similarly, NCLDR [feng2024ncldr] also uses nearest neighbors as signals, but its architecture is tailored to image data. In summary, existing SSL-integrated DR methods are either restricted to specific domains (e.g., image datasets) or designed for preprocessing rather than visualization. They do not provide a general solution for handling misaligned neighborhoods, diverse datasets, or visualization of high-dimensional data.
@@ III-A Motivation
High-dimensional data introduce fundamental challenges for DR methods with graph-based modeling, such as $t$-SNE [maaten2008visualizing], UMAP [mcinnes_umap_2020], and PaCMAP [wang2021understanding]. First, as dimensionality increases, conventional distance metrics (e.g., Euclidean and cosine distances) become more prone to the curse of dimensionality—pairwise distances measured in the high-dimensional space become more sensitive to noise within a data distribution [beyer1999nearest]. This makes it difficult to separate true neighbors from unrelated points when generating neighbor graphs from high-dimensional data. Second, many high-dimensional datasets (e.g., images or single-cell data in genomics) vary along curved low-dimensional manifolds defined by nonlinear generative factors such as pose or cell state. In these cases, close points in the high-dimensional space may be far apart along the underlying manifold and, therefore, may produce neighbor graphs where semantically distinct regions are connected. A recent DR method, LocalMAP [wang2025dimension], is designed to mitigate this issue in PaCMAP [wang2021understanding]. LocalMAP dynamically reweights neighbor edges and resamples neighbors during optimization. However, this approach relies on heuristic adjustments and does not resolve the broader issue of noisy variance and manifold-level distortions when constructing neighborhood graphs.
UMAP models the underlying manifold as a fuzzy graph derived from a kkNN graph that is constructed based on distances in high-dimensional space. Due to the above two-fold characteristics of high-dimensional data, the fuzzy graph can cause semantic mismatches, where points with no meaningful relationships are connected only because of distances in high-dimensional space rather than the underlying manifold geometry. As discussed in Sec. II-A, these semantic mismatches propagate into the final layout. This issue often manifests as loss of subcluster details in UMAP’s outcomes, when intra-cluster variation is not adequately represented, and as noisy or irregular cluster boundaries, when neighborhoods contain points belonging to different clusters. Therefore, with MAPLE, we aim to refine the graph construction phase of UMAP (see Phase 1 in Sec. II-A).
@@ III-C1 Step 1: Data Embedding
We apply this default architecture across datasets as a generic, lightweight feature extractor. For high-dimensional inputs ($D>512$), the first layer provides a moderate compression that helps filter out noise and stabilizes subsequent training. For lower-dimensional inputs ($D<512$), although the first layer could technically be reduced in size, we retain this expanded capacity to maintain a consistent and standardized architecture across datasets. Although more complex architectures such as ResNet [he2016deep] could be employed, we deliberately adopt this minimal encoder–projector design as an efficient and broadly applicable approach for high-dimensional data.
@@ III-C2 Step 2: Tentative Graph Construction
The neighbors in the tentative graph can be viewed as noisy “self-references”. We treat neighbors as similar data points and non-neighbors as dissimilar ones to enable self-supervised learning in Step 3. In the context of MVSSL, these neighbor data points can be regarded as augmented views of each data point, used to learn invariances. In our case, a data point and its neighbors act as multiple views reflecting the same local structure. However, unlike typical augmentation scenarios (e.g., creating augmented images via rotation), these views may not align perfectly with the same structure. Since the networks are initially untrained, early-epoch neighborhoods can still reflect noisy structure inherited from $X$. We therefore refine these neighborhoods through regularization with MMCRs.

### v2

@@ Abstract
We present a new nonlinear dimensionality reduction method, MAPLE, that enhances UMAP by improving manifold modeling. MAPLE employs a self-supervised learning approach to more efficiently encode low-dimensional manifold geometry. Central to this approach are maximum manifold capacity representations (MMCRs), which help untangle complex manifolds by compressing variances among locally similar data points while amplifying variance among dissimilar data points. This design is particularly effective for high-dimensional data with substantial intra-cluster variance and curved manifold structures, such as biological or image data. Our qualitative and quantitative evaluations demonstrate that MAPLE can produce clearer visual cluster separations and finer subcluster resolution than UMAP while maintaining a tractable computational cost.
@@ I Introduction
High-dimensional datasets are rapidly growing across scientific domains, from single-cell transcriptomics to large-scale image collections. Advances in genomics, machine learning, and related fields often produce measurements with tens of thousands of features, such as genes, pixels, or descriptors, that naturally place the data in very high dimensions. As humans cannot directly see high-dimensional spaces, effective visualization is essential to analyze these datasets. Dimensionality reduction (DR) methods make high-dimensional data visually explorable by projecting them into a low-dimensional space, typically 2D or 3D. Effective DR results can facilitate data annotation, reveal rare populations, and lead to the generation of new hypotheses.
@@
A wide range of DR methods has been proposed for visualization [espadoto2019toward, jeon2025unveiling]. Linear DR methods, including principal component analysis (PCA) [hotelling1933analysis], remain popular for their interpretability [cunningham2015linear, fujiwara2022ulca], but are limited in capturing complex, curved data structures. Nonlinear DR methods address this limitation by prioritizing local neighborhood relationships, allowing them to better represent data lying on complex manifolds [van2009dimensionality]. Among these, uniform manifold approximation and projection (UMAP) [mcinnes_umap_2020, healy2024umap] has become a method of choice across many domains, alongside alternatives such as $t$-SNE [maaten2008visualizing]. UMAP is valued for its scalability, graph-based formulation, and ability to capture nonlinear structure without strong parametric assumptions. UMAP constructs a weighted $k$-nearest neighbor (kkNN) graph in the input space to model an underlying manifold and then optimizes a low-dimensional layout so that the structure of the layout aligns with the graph according to a cross-entropy objective. This procedure is computationally efficient, and its cross-entropy objective often yields layouts with clear cluster separation.
@@
The quality of the resulting UMAP layout is strongly influenced by the quality of the weighted kkNN graph. In practice, this graph may include connections that bring semantically distinct points close together in the low-dimensional layout, and vice versa. We argue that a key limitation of UMAP lies not in the kkNN algorithm itself—which can serve as an effective proxy for nonlinear structure—but in how the graph is constructed, specifically the choice of data representation and distance space used to define neighborhood relationships. More specifically, standard kkNN construction based on raw high-dimensional data and predefined distance metrics faces the following two issues. First, as dimensionality increases, pairwise distances in high-dimensional space (e.g., Euclidean distances) tend to concentrate toward similar values [franccois2007concentration], making kkNN graphs increasingly sensitive to noise. Second, many datasets lie on curved manifolds [kocabas2019self], where points that appear as neighbors in the kkNN graph may actually belong to distinct semantic regions [conwell2024large]. In 2D or 3D layouts, these effects result in local clusters with spurious overlaps, thereby mixing groups that are semantically distinct. Recent theoretical work has also analyzed these issues of UMAP [damrich_umap_2021, bohm2022attraction], emphasizing its limitations in capturing subtle geometric differences.
@@
To address the above fundamental challenges, we introduce a novel nonlinear DR method, MAPLE (short for manifold-aware projection with learned edges). A central idea in MAPLE is that explicitly learning neighborhood relationships can substantially improve the quality of low-dimensional layouts. MAPLE refines neighborhood definitions by adapting multi-view self-supervised learning (MVSSL) [tsai2021selfsupervised, galvez2023role] to the context of nonlinear DR. Unlike standard methods that treat a weighted kkNN graph directly constructed from raw high-dimensional data as a fixed representation of the underlying manifold, MAPLE treats the initial neighbor relationships as weak, tentative self-supervised signals. We employ a learning phase to iteratively refine these signals. Instead of directly operating in the input high-dimensional space, we employ encoder and projector networks [grill2020bootstrap, kalantidis2021tldr, zbontar_barlow_2021] to map data into an embedding space where the distance relationships are learned rather than fixed by a predefined distance metric. A refined graph is then constructed based on these optimized distance relationships. The core of this learning phase is the maximum manifold capacity representation (MMCR) objective [yerxa2023mmcr], which compresses variance among locally similar points (i.e., neighbors) while amplifying variance across dissimilar ones (i.e., non-neighbors). The outcome of this learning phase is a refined, weighted kkNN graph that is subsequently laid out in a low-dimensional space using UMAP’s cross-entropy optimization.
@@
We evaluate MAPLE against multiple DR methods using a set of quality measures. Additionally, we qualitatively evaluate MAPLE through visual analysis on image classification datasets. We further demonstrate MAPLE’s utility in case studies on single-cell transcriptomics in collaboration with a bioinformatics expert (a co-author of this work). Our results show that MAPLE reveals fine-grained, meaningful patterns that support deeper downstream analysis.
Our contributions are as follows:
@@
Terminology. MAPLE produces two types of low-dimensional representations. To avoid ambiguity, we use the following terminology throughout this paper:
@@ II Background and Related Work
DR aims to extract low-dimensional representations from high-dimensional data while preserving meaningful structure. Nonlinear DR methods [van2009dimensionality] are designed to capture data patterns lying on nonlinear manifolds (e.g., curved surfaces). From the perspective of manifold modeling, nonlinear DR methods can be categorized by how they represent neighborhood structure. Graph-based modeling constructs a kkNN graph or a similar structure to capture local proximities along a manifold. This category includes early spectral and geodesic methods, such as ISOMAP [tenenbaum2000global] and Laplacian eigenmaps [belkin2003laplacian], as well as more recent methods, including LargeVis [tang2016visualizing], $t$-SNE [maaten2008visualizing], and UMAP [mcinnes_umap_2020]. Diffusion-based modeling (e.g., spectral embeddings [meila2001spectral] and PHATE [moon_PHATE_2019]) simulates diffusion processes (e.g., random walks) on the data and uses transition probabilities as similarities to preserve manifold structure. Kernel-based modeling [mika1998kernel, gisbrecht2015parametric] employs kernel functions to implicitly map data into higher-dimensional feature spaces, enabling the modeling of complex shapes. Our new method, MAPLE, belongs to the graph-based family. We specifically aim to advance graph-based modeling, using UMAP (a widely used state-of-the-art method) as our primary baseline.
@@ II-A UMAP
where $\rho_{i}$ and $\sigma_{i}$ are local connectivity parameters automatically determined by UMAP to distribute data points uniformly over the manifold. The weight, $w_{ij}$, reflects the strength of neighborhood relationships between points $i$ and $j$ (precisely, the membership strength of a 1-simplex between them [mcinnes_umap_2020]).
@@
$\arg\min_{\mathbf{Y}}\sum_{i\neq j}\Big[w_{ij}\log\!\left(\frac{w_{ij}}{v_{ij}}\right)+(1-w_{ij})\log\!\left(\frac{1-w_{ij}}{1-v_{ij}}\right)\Big].$ (2)
@@
As described in Eq. 1, UMAP assumes that data points are uniformly distributed on the manifold under some Riemannian metric. Under this assumption, any ball of fixed volume should contain approximately the same number of points, regardless of its position on the manifold. This motivates the use of fixed-size kkNNs to represent local relationships. In practice, however, real data are rarely uniform: they often vary in density or lie on curved manifolds. As a result, kkNNs constructed with a conventional metric (e.g., Euclidean distance) and the connectivity parameters, $\rho_{i}$ and $\sigma_{i}$, derived from them can misrepresent geodesic proximity. Since UMAP’s cross-entropy objective optimizes only over the weighted kkNN graph, any mismatch introduced during graph construction directly propagates into the final layout. This observation highlights a research gap: improving neighborhood construction is essential for achieving low-dimensional layouts that faithfully reflect the underlying manifold geometry.
@@ II-C Self-supervised Learning and Integration into DR
Integration of SSL into DR.  Several works have integrated SSL into DR to improve layout quality. These works share critiques of the graph construction phase, including unreliable distances in high-dimensional space [bohm2022unsupervised], semantically unstable neighborhoods [hu2023your], and distortion caused by curvature or density differences [kalantidis2021tldr, feng2024ncldr]. SSL objectives can help learn more semantically meaningful results. For example, $t$-SimCNE [bohm2022unsupervised] combines the contrastive learning method, SimCLR, with a $t$-SNE-like loss to improve similarity estimation of data points. However, as it relies on SimCLR, $t$-SimCNE is only applicable to image datasets. TLDR [kalantidis2021tldr] employs a single nearest neighbor as the self-supervised signal and Barlow Twins [zbontar_barlow_2021] as the SSL loss. While effective for data compression and preprocessing, TLDR is not designed for visualization purposes. Moreover, its reliance on a fixed $1$-nearest neighbor graph prevents it from addressing misaligned neighborhood issues. Similarly, NCLDR [feng2024ncldr] also uses nearest neighbors as signals, but its architecture is tailored to images.
In summary, existing SSL-integrated DR methods are either restricted to specific domains (e.g., image datasets) or designed for preprocessing rather than visualization. They do not provide a general solution for handling misaligned neighborhoods, diverse datasets, or visualization of high-dimensional data.
@@ III-A Motivation
High-dimensional data introduce fundamental challenges for DR methods with graph-based modeling, such as $t$-SNE [maaten2008visualizing], UMAP [mcinnes_umap_2020], and PaCMAP [wang2021understanding]. First, as dimensionality increases, conventional distance metrics (e.g., Euclidean and cosine distances) become more prone to the curse of dimensionality—pairwise distances measured in the high-dimensional space become more sensitive to noise within a data distribution [beyer1999nearest]. This makes it difficult to distinguish true neighbors from unrelated points when generating neighbor graphs from high-dimensional data. Second, many high-dimensional datasets (e.g., images or single-cell data in genomics) vary along curved low-dimensional manifolds defined by nonlinear generative factors such as pose or cell state. In these cases, close points in the high-dimensional space may be far apart along the underlying manifold and, therefore, may produce neighbor graphs where semantically distinct regions are connected. A recent DR method, LocalMAP [wang2025dimension], is designed to mitigate this issue in PaCMAP [wang2021understanding]. LocalMAP dynamically reweights neighbor edges and resamples neighbors during optimization. However, this approach relies on heuristic adjustments and does not resolve the broader issue of noisy variance and manifold-level distortions when constructing neighborhood graphs.
UMAP models the underlying manifold as a fuzzy graph derived from a kkNN graph that is constructed based on distances in high-dimensional space. Due to the above two-fold characteristics of high-dimensional data, the fuzzy graph can cause semantic mismatches, where points with no meaningful relationships are connected only because of distances in high-dimensional space rather than the underlying manifold geometry. As discussed in Sec. II-A, these semantic mismatches propagate into the final layout. This issue often manifests as loss of subcluster details in UMAP’s outcomes when intra-cluster variation is not adequately represented, and as noisy or irregular cluster boundaries when neighborhoods contain points belonging to different clusters. Therefore, with MAPLE, we aim to refine the graph construction phase of UMAP (see Phase 1 in Sec. II-A).
@@ III-C1 Step 1: Data Embedding
We apply this default architecture across datasets as a generic, lightweight feature extractor. For high-dimensional inputs ($D>512$), the first layer provides a moderate compression that helps filter out noise and stabilizes subsequent training. For lower-dimensional inputs ($D<512$), although the first layer could technically be reduced in size, we retain this expanded capacity to maintain a consistent and standardized architecture across datasets. Alternatively, the architecture could be scaled (e.g., linearly) with the input dimensionality; exploring such scaling strategies is a worthwhile direction for future work. Although more complex architectures such as ResNet [he2016deep] could be employed, we deliberately adopt this minimal encoder–projector design as an efficient and broadly applicable approach for high-dimensional data.
@@ III-C2 Step 2: Tentative Graph Construction
The neighbors in the tentative graph can be viewed as noisy “self-references”. We treat neighbors as similar data points and non-neighbors as dissimilar ones to enable self-supervised learning in Step 3. In the context of MVSSL, these neighbor data points can be regarded as augmented views of each data point, used to learn invariances. In our case, a data point and its neighbors act as multiple views reflecting the same local structure. However, unlike typical augmentation scenarios (e.g., creating augmented images via rotation), these views may not align perfectly with the same structure. Since the networks are initially untrained, early-epoch neighborhoods can still reflect noisy structure inherited from $\mathbf{X}$. We therefore refine these neighborhoods through regularization with MMCRs.